OpenStudy (anonymous):

Am I correct?? Use log5 3 = 0.683 and log5 6 = 1.113 to evaluate the logarithm 10. log5 9 Log 6 + Log 3 1.113 + 0.683 = 1.796

zepdrix (zepdrix):

log5(9) = log5(3*3) use your multiplication rule! :)

OpenStudy (anonymous):

What do i do after??

zepdrix (zepdrix):

log5(3*3) = log5(3) + log5(3) then use the information they gave you to replace your log5(3)'s

OpenStudy (anonymous):

so 0.683 + 0.683?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

yessss :)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

ya correct